A rectangle has a perimeter of 80 centimeter. Its length is 10 centimeter greater than its width. What is the area, in square centimeter, of the rectangle?
Answer by stanbon(75887)

A rectangle has a perimeter of 80 centimeter. Its length is 10 centimeter greater than its width. What is the area, in square centimeter, of the rectangle?
-----------------------
P = 2(L + W)
80 = 2(W+10 + W)
40 = 2W+10
2W = 30
W = 15
L = 15+10 = 25
-------------------------
Area = 15*25 = 375 cm^2
===================================
Cheers,
Stan H.